• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name yez

    Log entry time 05:50:59 on May 8,2011

    Entry number 358566

    keyword=High DAQ rate fixed

    Alexandre and I made a quick access to check the ECL cables connection on the trigger supervisor at Bigbite weldment. Basically what we did was to push all ECL connectors that connect to the trigger supervisor and make sure the high rate disppears with Cosmic Ray. As Alexandre said, this problem happened before durling SRC experiment, and at that time it was because of a bad connector of one ECL cable. I did not checked whether it was till because of the same bad cable, but we may still need to replace the cable this afternoon durling experiments switching. So far, the DAQ works normally


    FIGURE 1

    FIGURE 2